# Hyperparameter Optimization for Randomized Algorithms: A Case Study for Random Features

### v1

@@ 1 Introduction
Nonconvex optimization is ubiquitous in machine learning. In deep learning, it most frequently arises when training a neural network. This involves adjusting the weights and biases of the network to minimize a cost function. Alternatively, one can greatly simplify this optimization task by replacing the inner network weights with fixed samples from a probability distribution (Scardapane and Wang, 2017) and only learning the neural network’s linear output layer. With a suitable choice of loss function, such as a (possibly regularized) least squares functional, this optimization problem becomes a convex program. In the least squares example, its explicit solution involves the inversion of a linear system. Such an approach reduces approximation power, but does not have to deal with the challenges that come with nonconvexity. A large body of empirical evidence demonstrates that these convex reductions display remarkable performance in streaming large data sets at moderate processing and memory costs during training, partly because they do not require backpropagation (Chattopadhyay et al., 2020; Le et al., 2013; Sun et al., 2015). Performance on (often high-dimensional) problems with little data available for training is less explored (Gauthier et al., 2021b; Lanthaler and Nelsen, 2023).
@@
Motivated by scientific and engineering contexts, in this paper we instantiate the preceding systematic optimization pipeline for the specific machine learning task of function emulation from a small noisy data set. We apply the framework to one randomized algorithm, the method of random features (RFs, Rahimi and Recht, 2007). We benchmark the RF approach with the current state-of-the-art tool for such tasks, the kernel-based method of Gaussian processes (GPs, Williams and Rasmussen, 2006). The underlying hyperparameter training problem for the RF method involves optimally adapting the RF sampling distribution $\mu_{u}$ to data. This is equivalent to learning a data-adapted kernel or Bayesian prior distribution. Indeed, RFs, which in special cases can be viewed as neural networks with randomly sampled weights, by construction can also be viewed as low-rank approximations to GPs. A broad chronology of the developments and performance gains from RF approximation of kernel and GP methods can be found in the work of Liu et al. (2022).
@@ 2.1.1 Gaussian Process Regression
where the posterior mean function and posterior covariance function are given by
@@ 2.1.2 Random Feature Regression
for $x\in\mathcal{X}$ and $x^{\prime}\in\mathcal{X}$ and some $M\in\mathbb{N}$. This is a random low-rank prior covariance function. RFR is then simply the process of applying Bayes’ rule to the RF prior $\mathsf{GP}(\overline{F},K_{M})$ under the likelihood model (1) to obtain the RF posterior $\mathsf{GP}(\overline{F}_{M}^{(N)},K_{M}^{(N)})$. Here, the RF posterior mean $\overline{F}_{M}^{(N)}$ and RF posterior covariance $K_{M}^{(N)}$ are given by (6) except with all instances of $K$ replaced by $K_{M}$. In terms of sample paths, we write $F_{M}\sim\mathsf{GP}(\overline{F},K_{M})$ for the RF prior and $F_{M}\,|\,D_{N}\sim\mathsf{GP}(\overline{F}_{M}^{(N)},K_{M}^{(N)})$ for the RF posterior. Although it appears that RFR is just GPR with a special choice of prior, from an implementation and practical point of view it is provably beneficial to view RFR as an entirely distinct methodology (Rahimi and Recht, 2008a, b). Indeed, RFR can be interpreted as a finite rank Bayesian linear model in weight space (Bishop and Nasrabadi, 2006, Section 3.3), which simplifies sampling of the posterior distribution. Moreover, in what follows we show that RFR involves different system matrices that are cheaper to invert and coefficients that have different interpretations than those in appearing in GPR. We leave a rigorous analysis of the convergence of RFR to GPR to future work.
@@ 2.2.1 Vector-Valued Gaussian Process Regression
Although (20) visually mimics its scalar-valued counterpart (6), the computational effort is greatly magnified in the vector-valued setting. Effectively, the sample size $N$ is replaced by $Np$ in all scalar-valued output space GPR complexity estimates. Indeed, (20) requires solving $Np$ by $Np$ linear systems of equations instead of $N$ by $N$ ones. Table 1 summarizes the space, offline, and online costs of the approach. This poor scaling with output space dimension $p$ has severely limited the practical utility of vector-valued GPR to date. Indeed, existing studies usually only work with separable (or even scalar) matrix-valued kernels, which suffer from a simplistic rank-one tensor structure (Kadri et al., 2016; Owhadi, 2022), or diagonal kernels, which ignore correlations in the output space (Cleary et al., 2021). Vector-valued RFs enable scalable approximate GP inference without neglecting correlations in the output space.
@@ 3.1 Empirical Bayes Motivation
To overcome this, the empirical Bayes methodology takes an optimization approach (Chen et al., 2021). Assuming that $u$ is finite-dimensional and the prior on $u$ is uninformative, such as a uniform distribution, empirical Bayes maximizes the probability density of $u\,|\,Y$, which is the joint posterior $(F,u)\,|\,Y$ marginalized over $F$. By Bayes’ rule, the density of $u\,|\,Y$ is proportional to the marginal likelihood $Y\,|\,u$. Under our Gaussian data likelihood (19), $Y\,|\,u\sim\mathsf{N}(0,K^{(u)}(X,X)+B_{\Sigma})$ so that empirical Bayes minimizes the negative log marginal likelihood
@@ 4.3 Accelerated Uncertainty Quantification for a Cloud Parametrization Scheme
One approach to address the challenges of using MCMC is to replace the parameter-to-data (forward) map with a smooth and fast-evaluating surrogate (i.e., a statistical machine learning emulator) and sample using the surrogate instead. For example, when the forward map is random, authors of Cleary et al. (2021); Dunbar et al. (2021) approximated samples of the forward map by a trained Gaussian process $G\sim\mathsf{GP}(G_{\mathrm{ML}},\Gamma_{\mathrm{ML}})$. Then, (51) can be replaced with
@@ 5 Conclusion
When trained, random features can offer approximation with fewer assumptions and computational requirements than Gaussian processes. As illustrated in Section 4.3, vector-valued random features were able to produce qualitatively similar posterior distributions without relying on diagonalization, though the framework can take advantage of this when known. As with Gaussian processes, the choice of kernel or feature map is important. In this investigation, the random Fourier features parametrized by low-rank perturbations (subection 3.4) were effective across all test cases, given a suitably large rank. On the other hand, multi-output kernel structures that separate input and output spaces led to poor results. Further investigation of sensible kernel selection is left to future work.
@@
A fruitful area for future development for this case study is in alternative choices of loss function. More rigorous treatment of this approach may improve optimizer performance and accelerate convergence. Alternative loss functions not motivated from Bayesian arguments also exist for kernel methods (Owhadi and Yoo, 2019); such loss functions have proven successful in derivative-based settings for deterministic kernels and could also be recast for random objective functions in the present derivative-free framework. Furthermore, as EKI solves an inverse problem, misfits appearing in the cost function need not be directly supervised, and so this opens the door to training hyperparameters from indirect data. Indirect data is important for many applications where one embeds a machine learning surrogate in place of parametrization or model error compensation, and requires online training (e.g., Pahlavan et al., 2024; Wu et al., 2024).
More broadly, this work intends to act as a case study for a general framework for learning distributions that are sampled in randomized algorithms. Nearby examples are other machine learning tools that follow a randomization approach (Scardapane and Wang, 2017), such as Nyström methods (Williams and Seeger, 2000; Sun et al., 2015) or reservoir computers (Gauthier et al., 2021a). Common patterns to tune such methods often follow the literature of random features: employing manual calibration, fitting to one realization of randomness, and bespoke algorithms to solve the problem for a narrow choice of algorithm parameters. Our approach lays groundwork for new automated approaches of calibration which could accelerate research, as added robustness may lead to quicker uptake of such approaches.
@@
ORAD is supported by the generosity of Eric and Wendy Schmidt by recommendation of the Schmidt Futures program, the National Science Foundation (NSF) under award number AGS-1835860, the Defense Advanced Research Projects Agency (Agreement No. HR00112290030), the Heising–Simons Foundation, Audi Environmental Foundation, and the Cisco Foundation. NHN acknowledges support from the NSF Graduate Research Fellowship Program under award number DGE-1745301, the Amazon/Caltech AI4Science Fellowship, the Air Force Office of Scientific Research under MURI award number FA9550-20-1-0358 (Machine Learning and Physics-Based Modeling and Simulation), and the Department of Defense Vannevar Bush Faculty Fellowship held by Andrew M. Stuart under Office of Naval Research award number N00014-22-1-2790.

### v2

@@ 1 Introduction
Nonconvex optimization is ubiquitous in machine learning. In deep learning, it most frequently arises when training a neural network. This involves adjusting the weights and biases of the network to minimize a cost function. Alternatively, one can greatly simplify this optimization task by replacing the inner network weights with fixed samples from a probability distribution (Huang et al., 2006; Scardapane and Wang, 2017) and only learning the neural network’s linear output layer. With a suitable choice of loss function, such as a (possibly regularized) least squares functional, this optimization problem becomes a convex program. In the least squares example, its explicit solution involves the inversion of a linear system. Such an approach reduces approximation power, but does not have to deal with the challenges that come with nonconvexity. A large body of empirical evidence demonstrates that these convex reductions display remarkable performance in streaming large data sets at moderate processing and memory costs during training, partly because they do not require backpropagation (Chattopadhyay et al., 2020; Le et al., 2013; Sun et al., 2015; Bollt, 2021). Performance on (often high-dimensional) problems with little data available for training is less explored (Gauthier et al., 2021; Lanthaler and Nelsen, 2023).
@@
Motivated by scientific and engineering contexts, in this paper we instantiate the preceding systematic optimization pipeline for the specific machine learning task of function emulation from a small noisy data set. We apply the framework to one randomized algorithm, the method of random features (RFs, Rahimi and Recht, 2007). We benchmark the RF approach with the current state-of-the-art tool for such tasks, the kernel-based method of Gaussian processes (GPs, Williams and Rasmussen, 2006). The underlying hyperparameter training problem for the RF method involves optimally adapting the RF sampling distribution $\mu_{u}$ to data. This is equivalent to learning a data-adapted kernel or Bayesian prior distribution. Indeed, RFs, which in special cases can be viewed as neural networks with randomly sampled weights (e.g., “extreme learning machines” from Huang et al. (2006)), by construction can also be viewed as low-rank approximations to GPs. A broad chronology of the developments and performance gains from RF approximation of kernel and GP methods can be found in the work of Liu et al. (2022).
@@ 2.1.1 Gaussian Process Regression
where the posterior mean function and posterior covariance function are given by [^3] The posterior mean formula here is simply kernel ridge regression (Kanagawa et al., 2018) with a specific ridge penalty parameter depending on $\sigma^{2}$ and $N$.
@@ 2.1.2 Random Feature Regression
for $x\in\mathcal{X}$ and $x^{\prime}\in\mathcal{X}$ and some $M\in\mathbb{N}$. This is a random finite rank prior covariance function. RFR is then simply the process of applying Bayes’ rule to the RF prior $\mathsf{GP}(\overline{F},K_{M})$ under the likelihood model (1) to obtain the RF posterior $\mathsf{GP}(\overline{F}_{M}^{(N)},K_{M}^{(N)})$. Here, the RF posterior mean $\overline{F}_{M}^{(N)}$ and RF posterior covariance $K_{M}^{(N)}$ are given by (6) except with all instances of $K$ replaced by $K_{M}$. In terms of sample paths, we write $F_{M}\sim\mathsf{GP}(\overline{F},K_{M})$ for the RF prior and $F_{M}\,|\,D_{N}\sim\mathsf{GP}(\overline{F}_{M}^{(N)},K_{M}^{(N)})$ for the RF posterior. Although it appears that RFR is just GPR with a special choice of prior, from an implementation and practical point of view it is provably beneficial to view RFR as an entirely distinct methodology (Rahimi and Recht, 2008a, b). Indeed, RFR can be interpreted as a finite rank Bayesian linear model in weight space (Bishop and Nasrabadi, 2006, Section 3.3), which simplifies sampling of the posterior distribution. Moreover, in what follows we show that RFR involves different system matrices that are cheaper to invert and coefficients that have different interpretations than those in appearing in GPR. We leave a rigorous analysis of the convergence of RFR to GPR to future work.
@@ 2.2.1 Vector-Valued Gaussian Process Regression
Although (20) visually mimics its scalar-valued counterpart (6), the computational effort is greatly magnified in the vector-valued setting. Effectively, the sample size $N$ is replaced by $Np$ in all scalar-valued output space GPR complexity estimates. Indeed, (20) requires solving $Np$ by $Np$ linear systems of equations instead of $N$ by $N$ ones. Table 1 summarizes upper bounds on the space, offline, and online costs of the approach. This poor scaling with output space dimension $p$ has severely limited the practical utility of vector-valued GPR to date. Indeed, existing studies usually only work with separable (or even scalar) matrix-valued kernels, which suffer from a simplistic rank-one tensor structure (Kadri et al., 2016; Owhadi, 2022), or diagonal kernels, which ignore correlations in the output space (Cleary et al., 2021). Vector-valued RFs enable scalable approximate GP inference without neglecting correlations in the output space.
@@ 3.1 Empirical Bayes Motivation
To overcome this, the empirical Bayes methodology takes an optimization approach (Chen et al., 2021; Naslidnyk et al., 2023). Assuming that $u$ is finite-dimensional and the prior on $u$ is uninformative, such as a uniform distribution, empirical Bayes maximizes the probability density of $u\,|\,Y$, which is the joint posterior $(F,u)\,|\,Y$ marginalized over $F$. By Bayes’ rule, the density of $u\,|\,Y$ is proportional to the marginal likelihood $Y\,|\,u$. Under our Gaussian data likelihood (19), $Y\,|\,u\sim\mathsf{N}(0,K^{(u)}(X,X)+B_{\Sigma})$ so that empirical Bayes minimizes the negative log marginal likelihood
@@ 4.3 Accelerated Uncertainty Quantification for a Cloud Parametrization Scheme
One approach to address the challenges of using MCMC is to replace the parameter-to-data (forward) map with a smooth and fast-evaluating surrogate (i.e., a statistical machine learning emulator) and sample using the surrogate instead. For example, when the forward map is random, previous work (Cleary et al., 2021; Dunbar et al., 2021) approximated samples of the forward map by a trained Gaussian process $G\sim\mathsf{GP}(G_{\mathrm{ML}},\Gamma_{\mathrm{ML}})$. Then, (51) can be replaced with
@@ 5 Conclusion
When trained, random features can offer approximation with fewer assumptions and computational requirements than Gaussian processes. As illustrated in Section 4.3, vector-valued random features were able to produce qualitatively similar posterior distributions without relying on diagonalization, though the framework can take advantage of this when known. As with Gaussian processes, the choice of kernel or feature map is important. In this investigation, the random Fourier features parametrized by low-rank perturbations (Subsection 3.4) were effective across all test cases, given a suitably large rank. On the other hand, multi-output kernel structures that separate input and output spaces led to poor results. Further investigation of sensible kernel selection is left to future work.
@@
A fruitful area for future development for this case study is in alternative choices of loss function. More rigorous treatment of this approach may improve optimizer performance and accelerate convergence. Alternative loss functions not motivated from Bayesian arguments also exist for kernel methods and are based on cross validation (Owhadi and Yoo, 2019; Chen et al., 2021; Naslidnyk et al., 2023); such loss functions have proven successful in derivative-based settings for deterministic kernels and could also be recast for random objective functions in the present derivative-free framework. Furthermore, as EKI solves an inverse problem, misfits appearing in the cost function need not be directly supervised, and so this opens the door to training hyperparameters from indirect data. Indirect data is important for many applications where one embeds a machine learning surrogate in place of parametrization or model error compensation, and requires online training (e.g., Pahlavan et al., 2024; Wu et al., 2024).
More broadly, this work intends to act as a case study for a general framework for learning distributions that are sampled in randomized algorithms. Nearby examples are other machine learning tools that follow a randomization approach (Scardapane and Wang, 2017), such as Nyström methods (Williams and Seeger, 2000; Sun et al., 2015; Meanti et al., 2022) or reservoir computers (Gauthier, 2018; Gauthier et al., 2021; Griffith et al., 2019). Common patterns to tune such methods often follow the literature of random features: employing manual calibration, fitting to one realization of randomness, and bespoke algorithms to solve the problem for a narrow choice of algorithm parameters. Recent work in these fields has seen other Bayesian optimization methods learn low-dimensional (fewer than $10$) hyperparameters effectively (e.g., Griffith et al. (2019); Racca and Magri (2021); Antonik et al. (2023)). As EKI is highly scalable when compared with such Gaussian process-based optimizers, it may enable the learning of much higher dimensional hyperparameters. Our approach lays groundwork for new automated approaches of calibration which could accelerate research, as added robustness may lead to quicker uptake of such approaches.
@@
ORAD is supported by Schmidt Sciences, LLC, the National Science Foundation (NSF) under award number AGS-1835860, the Defense Advanced Research Projects Agency (Agreement No. HR00112290030). NHN acknowledges support from the NSF Graduate Research Fellowship Program under award number DGE-1745301, the Amazon/Caltech AI4Science Fellowship, the Air Force Office of Scientific Research under MURI award number FA9550-20-1-0358 (Machine Learning and Physics-Based Modeling and Simulation), and the Department of Defense Vannevar Bush Faculty Fellowship held by Andrew M. Stuart under Office of Naval Research award number N00014-22-1-2790. The authors are grateful to Daniel J. Gauthier for helpful comments on a previous version of the paper.